### v1

# cuPDLP+: A Further Enhanced GPU-Based First-Order Solver for Linear Programming

Abstract
We introduce cuPDLP+, a further enhanced GPU-based first-order solver for linear programming. Building on the predecessor cuPDLP, cuPDLP+ incorporates recent algorithmic advances, including the restarted Halpern PDHG method with reflection, a novel restart criterion, and a PID-controlled primal weight update. These innovations are carefully tailored for GPU architectures and deliver substantial empirical gains. On a comprehensive benchmark of MIPLIB LP relaxations, cuPDLP+ achieves 2x – 4x speedup over cuPDLP, with particularly strong improvements in high-accuracy and presolve-enabled settings.
1 Introduction
Linear programming (LP) is a central class of mathematical optimization problems due to its broad applicability, mathematical elegance, and computational efficiency [4]. The two classic algorithmic frameworks for solving LPs are the simplex method [8, 7] and interior-point methods (IPMs) [10, 18], both of which have shaped decades of theoretical development and practical deployment. The simplex method, introduced by Dantzig in the 1940s [8], proceeds by traversing the vertices of the feasible polyhedron and has demonstrated remarkable efficiency in practice despite its exponential worst-case complexity. Its numerical robustness and interpretability have made it a mainstay in commercial solvers. In contrast, IPMs operate within the interior of the feasible region, iteratively following a central path toward optimality. Since the seminal work of Karmarkar in the 1980s [10], IPMs have offered polynomial-time theoretical guarantees and have been extended to handle large and structured problems with high numerical precision. Modern LP solvers often integrate both approaches, exploiting the simplex method’s flexibility and warm-start capability alongside the theoretical convergence guarantees and scalability of IPMs. These methods remain the foundation of the state-of-the-art CPU-based LP solvers, underpinning their ability to deliver high-accuracy solutions across diverse problem instances.
Recently, first-order methods (FOMs) [16] have emerged as a compelling alternative for solving large-scale LPs, offering a new paradigm that complements the traditional dominance of simplex and interior-point methods. Unlike simplex and interior-point methods, which rely on matrix factorizations and sequential computations, FOMs operate using simple iterative updates, primarily matrix-vector multiplications, making them particularly attractive for large-scale LPs. A notable breakthrough in this space is PDLP [1, 2], a solver based on the primal-dual hybrid gradient method (PDHG) [5] and tailored specifically for LP. By incorporating various practical enhancements, PDLP achieves significantly improved numerical stability and convergence in practice, positioning it as a practical and scalable solver for large-scale LPs within the FOM framework.
Building on the CPU-based PDLP, recent efforts have turned toward leveraging graphics processing units (GPUs) to further accelerate LP solvers. GPUs offer massive parallelism and high memory bandwidth, making them particularly well-suited for the core computational kernels of first-order methods, namely, sparse matrix-vector multiplications and vector operations. As LP problem sizes continue to grow, sometimes even reaching billions of variables and constraints in modern applications [15], harnessing GPU architectures becomes increasingly critical for achieving scalable performance of first-order methods. This shift has led to the development of a new generation of GPU-accelerated LP solvers that combine algorithmic efficiency with hardware-aware implementation strategies.
Particularly, cuPDLP [12, 14] is a GPU-accelerated solver that extends PDLP by adapting its first-order framework to modern GPU architectures, with several algorithmic changes to make it more suitable to GPU architecture. The initial implementation, developed in Julia and referred to as cuPDLP [12], offloads key linear algebra operations, such as sparse matrix-vector multiplications, to the GPU and incorporates GPU-friendly heuristics, achieving notable speedups on medium-to-large-scale LP instances. Serving as both a practical solver and a research platform, cuPDLP illustrates how careful alignment between algorithm design and hardware capabilities can yield substantial performance improvements. A subsequent C implementation, cuPDLP-C [14], was developed to facilitate integration with production-scale computing environments. These advances in GPU-based LP solvers have attracted strong interest from both optimization software companies and technology firms, and have influenced the design of several commercial solvers, including Gurobi, COPT, FICO Xpress and NVIDIA’s cuOpt.
More recently, a distinct approach based on the Halpern Peaceman–Rachford (HPR) method [17] was proposed. Building on this framework, HPR-LP [6] was developed as a GPU-based LP solver, and numerical results on standard benchmarks demonstrate superior performance compared to cuPDLP, especially in obtaining high-accuracy solutions.
These recent GPU-based solvers demonstrate the growing interest and practical potential of first-order methods when carefully adapted to modern hardware. Building on these foundations, we present cuPDLP+, a further enhanced GPU-accelerated linear programming solver. Compared to its predecessor cuPDLP, cuPDLP+ introduces several fundamental algorithmic changes that are primarily motivated by recent theoretical developments.
  • While cuPDLP is based on the restarted averaged PDHG (raPDHG) method [3], cuPDLP+ adopts the restarted Halpern PDHG (rHPDHG), a refinement inspired by recent theoretical insights[13]. This change allows the algorithm to take more aggressive steps, resulting in stronger empirical performance. The observed numerical improvements of cuPDLP+ over cuPDLP highlight the practical advantages of the rHPDHG framework.
  • In addition, cuPDLP+ explores a distinct set of enhancements and heuristics that further accelerate performance. These include a constant step-size rule, a restart condition aligned with the theoretical guarantees of rHPDHG [13], and a novel PID controller for updating the primal weight.
These algorithmic improvement translate into strong empirical performance of cuPDLP+. In various scenarios, cuPDLP+ achieves a 2× to 4× speedup. For example, in the high-accuracy setting with presolve, cuPDLP+ achieves a 2.9× speedup on 383 MIPLIB instances and a 4.63× speedup on hard instances, compared to cuPDLP.
1.1 Paper organization
Section 2 briefly introduces the form of LP to solve and the base algorithm vanilla PDHG. In Section 3, several algorithmic enhancement on top of cuPDLP are proposed to accelerate convergence. The numerical comparisons between cuPDLP+ and cuPDLP are presented in Section 4.
1.2 Notations
For a symmetric positive semidefinite matrix $M\in\mathbb{R}^{n\times n}$, the $M$-norm of a vector $x\in\mathbb{R}^{n}$ is defined as $\|x\|_{M}:=\sqrt{x^{\top}Mx}$. The orthogonal projector $\operatorname{proj}_{C}(x)$ denotes the Euclidean projection of a point $x\in\mathbb{R}^{n}$ onto a closed convex set $C\subseteq\mathbb{R}^{n}$, i.e., $\operatorname{proj}_{C}(x):=\arg\min_{y\in C}\|x-y\|_{2}$. For a set $C\subseteq\mathbb{R}^{n}$, we define $-C:=\{-x:x\in C\}$ as the reflection of $C$ about the origin. The $(i,j)$-th entry for a matrix $M\in\mathbb{R}^{m\times n}$ is denoted by $(M)_{ij}$. Given two vectors ({}$\ell\in(\mathbb{R}\cup\{-\infty\})^{n}$ and $u\in(\mathbb{R}\cup\{\infty\})^{n}$ satisfying $\ell\leq u$, define the mapping $p:\mathbb{R}^{n}\times(\mathbb{R}\cup\{-\infty\})^{n}\times(\mathbb{R}\cup\{\infty\})^{n}\to\mathbb{R}\cup\{\infty\}$ given by $p(y;\ell,u):=u^{\top}y^{+}-\ell^{\top}y^{-}$.
2 Preliminaries
In this section, we introduce the LP form that cuPDLP+ solves, followed by discussion on vanilla PDHG for solving LPs.
2.1 Linear programming
cuPDLP+ solves LP with the following primal-dual form:
$\min_{x\in\mathcal{X}}~{}~{}&~{}c^{\top}x&\\ \text{subject to:}~{}~{}&~{}Ax\in\mathcal{S}\ ,\end{aligned}\qquad\qquad\qquad\begin{aligned} \max_{y\in\mathcal{Y},r\in\mathcal{R}}\quad&-p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})\\ \text{subject to:}\quad&c-A^{\top}y=r\ ,\end{aligned}$ (1)
where $\mathcal{X}:=\{x\in{\mathbb{R}}^{n}:\ell_{v}\leq x\leq u_{v}\}$ with $\ell_{v}\in({\mathbb{R}}\cup\{-\infty\})^{n}$ and $u_{v}\in({\mathbb{R}}\cup\{\infty\})^{n}$, $\mathcal{S}:=\{s\in{\mathbb{R}}^{m}:\ell_{c}\leq s\leq u_{c}\}$ with $\ell_{c}\in({\mathbb{R}}\cup\{-\infty\})^{m}$ and $u_{c}\in({\mathbb{R}}\cup\{\infty\})^{m}$, $A\in{\mathbb{R}}^{m\times n}$, $c\in{\mathbb{R}}^{n}$, and the sets $\mathcal{Y}\subseteq{\mathbb{R}}^{m}$ and $\mathcal{R}\subseteq{\mathbb{R}}^{n}$ are Cartesian products with their $i$th components given by
$&\mathcal{Y}_{i}:=\begin{cases}\{0\}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{c})_{i}\in{\mathbb{R}},~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\quad\text{ and }\qquad\end{aligned}\begin{aligned} &\mathcal{R}_{i}:=\begin{cases}\{0\}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{v})_{i}\in{\mathbb{R}},~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\ .\end{aligned}$
Equivalently the primal-dual form of (1) is
$\max_{y\in\mathcal{Y}}\min_{x\in\mathcal{X}}\ L(y,x):=-p(y;-u_{c},-\ell_{c})+y^{\top}Ax+c^{\top}x\ .$ (2)
This LP form is used in CPU-based PDLP [2].
2.2 PDHG
PDHG serves as the base routine of cuPDLP+ to solve primal-dual problem (2). Specifically, the update rule of PDHG on (2) is given as:
$x^{k+1}=\operatorname{proj}_{\mathcal{X}}\left({x^{k}-{\tau}(c-A^{\top}y^{k})}\right)$ (3)
$y^{k+1}=y^{k}-\sigma A(2x^{k+1}-x^{k})-\sigma\operatorname{proj}_{-\mathcal{S}}\left({\sigma^{-1}y^{k}-A(2x^{k+1}-x^{k})}\right)\ ,$
where $\sigma$ and $\tau$ are dual and primal stepsizes respectively. cuPDLP+ further reparameterizes as $\tau=\frac{\eta}{\omega}$ and $\sigma=\eta\omega$, where $\eta$ is called the stepsize and $\omega$ is the primal weight. And the canonical norm of PDHG is defined by $\|\cdot\|_{P}$, where $P:=P_{\eta,\omega}=\begin{bmatrix}\frac{\omega}{\eta}I&A^{\top}\\ A&\frac{1}{\eta\omega}I\end{bmatrix}$. For notational convenience, we define the primal-dual iterate at iteration $k$ as $z^{k}=\begin{bmatrix}x^{k}\ y^{k}\end{bmatrix}$, and write $z^{k+1}=\operatorname{PDHG}(z^{k})$ to denote a single PDHG update step applied to $z^{k}$ as defined in (3).
3 Algorithmic enhancements on top of cuPDLP
To improve the efficiency and convergence behavior of cuPDLP+, we introduce a series of algorithmic enhancements on top of cuPDLP , inspired by recent advances in first-order methods [13]. These enhancements, implemented in our further enhanced solver cuPDLP+ , build on the PDHG framework and include techniques such as the Halpern iteration scheme, reflected updates, adaptive restarting, stepsize selection and dynamic primal-dual weight. Each component is carefully designed to accelerate convergence and enhance stability for solving linear programming problems. We detail these contributions below.
Halpern scheme. The Halpern scheme and its reflected variant are recent enhancements to the PDHG algorithm, aimed at improving convergence properties and accelerating solver performance. Originally developed to accelerate general operator splitting methods, the Halpern scheme has been successfully adapted to linear programming such as HPR-LP [6] and MPAX [11].
Halpern PDHG interpolates between the current PDHG iterate and the initial point, using a weighted average. Specifically, the update rule at iteration $k$ is:
$z^{k+1}=\frac{k+1}{k+2}\operatorname{PDHG}(z^{k})+\frac{1}{k+2}z^{0}\ .$
Reflection. The reflection technique builds upon the Halpern scheme by applying a reflected version of the PDHG operator, $(1+\gamma)\operatorname{PDHG}-\gamma\operatorname{id}$ with reflection $\gamma\in[0,1]$, instead of the vanilla PDHG update. This leads to the Reflected Halpern PDHG update:
$z^{k+1}=\frac{k+1}{k+2}\left({(1+\gamma)\operatorname{PDHG}(z^{k})-\gamma z^{k}}\right)+\frac{1}{k+2}z^{0}\ .$
The reflection mechanism effectively takes a longer step compared to the vanilla Halpern update and has been shown to improve convergence both in theory and in practice [13].
Adaptive restart. Restarting is a key enhancement on first-order methods for attaining high-accuracy solutions. A restart is triggered when certain progress criteria are met, and the algorithm restarts from a new initial solution. In the Halpern scheme, the anchor is periodically reset to the current solution. This helps the algorithm stay focused on the neighborhood of the optimal solution, especially as the initial anchor becomes increasingly outdated during iterations. Such restart strategies have been shown to accelerate convergence rates for Halpern PDHG [13].
cuPDLP+ adopts an adaptive restart strategy that evaluates potential restart conditions at each iteration. The core idea is to monitor a fixed-point error metric $r(z)=\|z-\operatorname{PDHG}(z)\|_{P}$ at solution $z$, and trigger a restart when specific decay patterns are observed. This fixed-point error metric is motivated by the recent theoretical insight [13], and it is different from the normalized duality gap used in CPU-based PDLP[1, 2] and the KKT error used in cuPDLP[12]. In addition, three restart conditions are used:
  • (sufficient decay) $r(z^{n,k})\leq\beta_{\mathrm{sufficient}}r(z^{n,0})$
  • (necessary decay + no local progress) $r(z^{n,k})\leq\beta_{\mathrm{necessary}}r(z^{n,0})$ and $r(z^{n,k})>r(z^{n,k-1})$
  • (artificial restart) $k\geq\beta_{\mathrm{artificial}}T$, where $T$ is the total iteration.
Stepsize. In contrast to an adaptive stepsize heuristic used in cuPDLP, cuPDLP+ switches to constant stepsize $\eta=\frac{0.99}{\|A\|_{2}}$, where operator norm of constraint matrix $\|A\|_{2}$ is approximated by power iteration.
Primal weight update. The primal weight $\omega$ is designed to balance the progress in the primal and dual spaces by adjusting its value. Specifically, the intuition is to determine the primal weight $\omega^{n}$ such that the weighted distances to optimality in the primal and dual domains are in the same scale, i.e., $\|(x^{n,t}-x^{*}),0\|_{w^{n}}\approx\|(0,y^{n,t}-y^{*})\|_{w^{n}}$. However, the optimal solution $(x^{,}y^{)}$ is not known during the iterations. To address this, cuPDLP estimates the primal and dual distances to optimality based on the observed movements in the previous epoch. The distance estimates are smoothed using exponential averaging and subsequently used to update the primal weight at each restart.
In cuPDLP+, we further enhance this strategy by modeling the update as a control problem and introducing a PID controller to dynamically regulate the primal weight. We define the error as the gap between the primal and dual distances on a logarithmic scale:
$e^{n}=\log\left(\frac{\sqrt{w^{n}}\left\|x^{n,t}-x^{*}\right\|_{2}}{\frac{1}{\sqrt{w^{n}}}\left\|y^{n,t}-y^{*}\right\|_{2}}\right)$
The primal weight is then updated according to:
$\log w^{n+1}=\log w^{n}-[K_{P}\cdot e^{n}+K_{I}\cdot\sum_{i=1}^{n}e^{i}+K_{D}\cdot(e^{n}-e^{n-1})]$
where $K_{P}$, $K_{I}$ and $K_{D}$ are the coefficients for the proportional, integral and derivative terms, respectively. The update is applied at each restart occurrence and the initial primal weight is set to $1.0$.
4 Numerical experiments
In this section, we compare the numerical performance of cuPDLP+ with cuPDLP. We first describes the setup of the experiments, followed by presents the numerical results on LP relaxations of instances from MIPLIB 2017 collection.
Benchmark datasets. We conduct numerical experiments on a widely used LP benchmark dataset: MIP Relaxations, which consists of 383 instances derived from root-node LP relaxations of MIPLIB 2017 [9]. The 383 instances in the MIP relaxations dataset are selected based on the same filtering criteria outlined in [12]. Based on the number of non-zeros of the constraint matrices, we further categorize these 383 instances into three groups, as summarized in Table 1.
Small Medium Large
Number of nonzeros 100K - 1M 1M - 10M >10M
Number of instances 269 94 20
Table 1: Scale and number of instances in MIP Relaxations.
Software. We implement cuPDLP+ in Julia, utilizing CUDA.jl as the interface for working with NVIDIA CUDA GPUs. We compare the performance of cuPDLP+ with cuPDLP. The running time of both cuPDLP+ and cuPDLP is measured after pre-compilation in Julia.
Computing environment. We use NVIDIA H100-SXM-80GB GPU, with CUDA 12.4. The experiments are performed in Julia 1.11.5.
Initialization. cuPDLP+ uses all-zero vectors as the initial starting points.
Optimality termination criteria. cuPDLP+ terminate when the relative KKT error is no greater than the termination tolerance $\epsilon\in(0,\infty)$:
$|c^{\top}x+p(-y;\ell_{c},u_{c})+p(-r;\ell_{v},u_{v})|\leq\epsilon(1+|p(-y;\ell_{c},u_{c})+p(-r;\ell_{v},u_{v})|+|c^{\top}x|)$
$\left\|Ax-\operatorname{proj}_{[L,U]}(Ax)\right\|_{2}\leq\epsilon\left({1+\left\|(L,U)\right\|_{2}}\right)$
$\left\|c-A^{\top}y-r\right\|_{2}\leq\epsilon(1+\|c\|_{2})$
$\left\|r-\operatorname{proj}_{\mathcal{R}}(r)\right\|_{2}\leq\epsilon(1+\|c\|_{2})\ .$
The termination criteria are checked for the original LP instance, not the preconditioned ones, so that the termination is not impacted by the preconditioning. In the experiment, we set $\epsilon=10^{-4}$ for moderate accuracy and $\epsilon=10^{-8}$ for high accuracy.
Time limit. We apply a time limit of 3600 seconds to small and medium-sized problems and 18000 seconds to large-scale problems.
Shifted geometric mean. We use the shifted geometric mean of solving time to evaluate solver performance across a collection of instances. Formally, the shifted geometric mean is defined as $\left(\prod_{i=1}^{n}(t_{i}+\Delta)\right)^{1/n}-\Delta$ where $t_{i}$ is the solve time for the $i$-th instance. We shift by $\Delta=10$ and denote this metric as SGM10. If an instance is unsolved, its solving time is set to the corresponding time limit.
Tables 2 and 3 compare the performance of cuPDLP and cuPDLP+ on 383 MIPLIB instances, with and without Gurobi presolve, respectively. Several key observations can be made:
  • Across all four settings, cuPDLP+ consistently outperforms cuPDLP in terms of overall numerical performance, solving more instances and achieving notable reductions in running time.
  • The performance gain of cuPDLP+ is especially significant for Small and Medium instances, while the improvement is less significant for Large instances, likely due to the high variance among the relatively small number of Large instances.
Table 4 reports the speedup of cuPDLP+ over cuPDLP. In particular, the first row focuses on hard instances, defined as those requiring at least 10 seconds to solve using either cuPDLP or cuPDLP+. In various scenarios, cuPDLP+ achieves a 2× to 4× speedup. The improvement is more substantial when Gurobi presolve is enabled compared to the case without presolve. The speedup is also more significant on harder instances. For example, in the high-accuracy setting with presolve, cuPDLP+ achieves a 2.9× speedup on all instances and a 4.63× speedup on hard instances.
Small (269)
(1-hour limit)
Medium (94)
(1-hour limit)
Large (20)
(5-hour limit)
Total (383)
Count Time Count Time Count Time Count Time
$\mathbf{10^{-4}}$ cuPDLP 266 8.81 91 11.55 19 77.43 376 11.07
[…]
$\mathbf{10^{-8}}$ cuPDLP 260 24.80 87 36.90 17 208.66 364 31.22
[…]
Table 2: Solve time in seconds and SGM10 on instances of MIP Relaxations without presolve.
Small (269)
(1-hour limit)
Medium (94)
(1-hour limit)
Large (20)
(5-hour limit)
Total (383)
Count Time Count Time Count Time Count Time
$\mathbf{10^{-4}}$ cuPDLP 268 5.11 92 9.03 19 26.74 379 6.75
[…]
$\mathbf{10^{-8}}$ cuPDLP 264 18.50 90 29.40 19 63.68 373 22.42
[…]
Table 3: Solve time in seconds and SGM10 on instances of MIP Relaxations with Gurobi presolve.
Without
Presolve
With
Presolve
1e-4 1e-8 1e-4 1e-8
Hard 3.30 3.35 4.05 4.63
Overall 2.01 2.28 2.49 2.90
Table 4: Summarization of cuPDLP+ speedup over cuPDLP. The first row reports the speedup of cuPDLP+ on instances where cuPDLP or cuPDLP+ takes more than 10 seconds to solve, and the second row summarizes the overall speedup across all instances.
Acknowledgement
Haihao Lu is supported by AFOSR Grant No. FA9550-24-1-0051 and ONR Grant No. N000142412735. Zedong Peng is supported by ONR Grant No. N000142412735. Jinwen Yang is supported by AFOSR Grant No. FA9550-24-1-0051.